# Take Your Best Shot: Sampling-Based Next-Best-View Planning for Autonomous Photography & Inspection

### v1

@@ Abstract
Autonomous mobile robots (AMRs) equipped with high-quality cameras have revolutionized the field of inspections by providing efficient and cost-effective means of conducting surveys. The use of autonomous inspection is becoming more widespread in a variety of contexts, yet it is still challenging to acquire the best inspection information autonomously. In situations where objects may block a robot’s view, it is necessary to use reasoning to determine the optimal points for collecting data. Although researchers have explored cloud-based applications to store inspection data, these applications may not operate optimally under network constraints, and parsing these datasets can be manually intensive. Instead, there is an emerging requirement for AMRs to autonomously capture the most informative views efficiently. To address this challenge, we present an autonomous Next-Best-View (NBV) framework that maximizes the inspection information while reducing the number of pictures needed during operations. The framework consists of a formalized evaluation metric using ray-tracing and Gaussian process interpolation to estimate information reward based on the current understanding of the partially-known environment. A derivative-free optimization (DFO) method is used to sample candidate views in the environment and identify the NBV point. The proposed approach’s effectiveness is shown by comparing it with existing methods and further validated through simulations and experiments with various vehicles.
Note—Code and videos of the simulations and experiments are provided in the supplementary material and can be accessed at https://www.bezzorobotics.com/sg-lb-iros24.
@@ I Introduction
Autonomous mobile robots (AMRs) are becoming increasingly common in a wide range of industries, including manufacturing, healthcare, logistics, and entertainment. Because of their agility and mobility, autonomous robots have been particularly useful for photography [1] and inspection purposes [2]. The deployment of robots for these applications also offers other benefits, including increased safety, efficiency, and accuracy. Robots equipped with cameras, depth sensors, and other tools can access areas that are hazardous or inaccessible to human inspectors [3] (e.g., disaster sites, high-rises, internal pipe systems, and underwater structures). The use of robots to inspect buildings and other infrastructures is a promising solution to address the challenges associated with traditional inspection methods. However, most of the existing autonomous inspection frameworks are not fully automated and collect all available information during operations, relegating data processing and decision-making to humans. Thus, to fully automate robotic inspection, robots must be able to assess the quality of inspection photos and discard unnecessary data. However, the evaluation of photos is inherently subjective. Even among human observers, reaching a unanimous agreement on the quality rating of a photo can be challenging [4, 5]. This introduces additional challenges in constructing a formal metric that precisely defines the criteria for a “good” picture. With these considerations in mind, we identify two notable gaps in the robotics literature focused on inspection and photography: i) the absence of an explicit definition of a good picture, and ii) that current inspection frameworks do not incentivize efficient data collection during prolonged tasks, overlooking the cost of consuming repetitive information and the need for extensive human post-processing efforts. We propose that bridging these gaps will make autonomous inspection more useful and accessible to real-world operations.
Figure 1: A pictorial depiction of the problem. A UAV is performing an inspection task to capture an oblique view (i.e., front view) of a house. When comparing these four examples, a bird’s eye view in D preserves the front view and provides the most comprehensive information.
Typically, during inspection tasks, the location and dimensions of the region of interest (ROI) are known, and the surroundings of the ROI are uncertain. For example, in the house inspection shown in Fig. 1, the location of the building is known, while the landscape and adjacent structures can change, potentially obstructing the ROI and jeopardizing the safety of the unmanned aerial vehicle (UAV). In this scenario, the UAV should autonomously detect that trees obstruct viewing point (A) and that, by viewing from either side (B, C), more information can be obtained. The UAV’s objective is to autonomously evaluate different viewpoint candidates without visiting them, identify the best viewpoint (a bird’s-eye view, D in the example in Fig. 1), and route to a new viewpoint, considering the observed obstacles.
@@
We consider the framework presented in this paper as a Next-Best-View (NBV) solution in which the goal is to obtain the fewest viewing points and maximize inspection information. We propose an evaluation metric based on the perspective distortion, the scale of a target within the viewing frame, and the estimated target coverage which will allow us to find the best viewpoint. The metric is used in conjunction with a derivative-free optimization (DFO) method which samples the environment to find the best inspection point. The main contribution of this work is a computationally efficient Gaussian process interpolation using DFO to optimize the proposed evaluation metric over uniformly sampled candidate views maximizing target coverage at runtime.
@@ II Related Work
A well-studied problem in the field of mobile robot inspection is the Next-Best-View problem. First introduced by Connolly [6], NBV focuses on developing algorithms to efficiently plan and execute the exploration of objects or environments. The main goal of NBV is to select the most informative viewpoint for the robot to observe and improve its knowledge of a mission objective, such as an environment or an object. This is achieved by computing an optimized viewpoint based on a task-oriented optimization term that considers the robot’s objectives and constraints. As NBV focuses on determining the optimal viewpoint for the next observation, it can be considered as a subset or specific case within the exploration framework. In the literature, NBV has been extensively explored for 3D object reconstruction. One of the challenges in these applications is how to choose the NBV. Authors in [7] employ a weighted multi-objective optimization approach to select NBVs for a mobile robotic arm, while [8] uses a double branch network architecture to rank NBVs. Although many applications involve a single robot, [9] leverages two robotic arms and multi-agent systems, respectively, to select NBVs that promote efficiency and achieve faster reconstruction. Although these works demonstrate the effectiveness of NBV in object reconstruction, they mainly focus on dense data collection, encouraging robots to gather data close to the target without considering repetitive data collection or the overall aesthetics of the target.
Another challenge of applying the NBV problem is computing the optimal viewpoint in partially known or unknown environments, where unexpected objects can not only obscure the target but also pose safety threats to vehicles. Although some research focuses on navigating robots to optimal viewpoints in known environments [10], several works address this challenge in conjunction with NBV. Authors in [11] investigate a receding-horizon NBV, utilizing a random tree method to guide the robot along a path in an unknown environment. [12] proposes a guided NBV approach for large-scale 3D reconstruction, which requires a rough global scan prior to a detailed NBV inspection. [13] computes NBV based on the map built during frontier exploration. Other works related to the NBV focus on exploring unknown environments [14]. In our work, the robot starts with the target’s location as a priori information and dynamically updates the NBV to efficiently adapt to changes in the environment, ensuring effective guidance of the vehicle.
@@ III Problem Formulation
For inspection tasks, it is usually the case that the target information is known (i.e., the location of a house, pipeline, or structure) while the surroundings can be uncertain. In this work, we assume that the robot knows the target information, $\mathbf{T}$, including the location and dimensions. Obstacles in the environment, denoted as $\mathcal{O}$, can either be known or unknown and are represented by an occupancy map $\mathcal{M}$. To find the best point for inspection, the vehicle should update the best view position as obstacles are detected that compromise inspection photo quality. We assume that the vehicle is equipped with a range sensor (i.e., LiDAR, sonar, cameras) which enables the robot to recognize obstacles and measure its distance to the target within the vehicle’s field of view. The goal is to maximize the inspection information while reducing the number of pictures needed during operations. To achieve this, we formally define the problems as follows:
@@
Metrics of the Best View for Target Inspection: Consider the location and dimensions of the target represented by the coordinate set $\mathbf{T}$ and a set of initially unknown obstacles that are discovered at runtime and represented by an occupancy map $\mathcal{M}$. The goal is to find an evaluation metric $G(\mathbf{T},\mathcal{M},\mathbf{P})$ that scores the sampled viewpoints $\mathbf{P}$ for the vehicle to visit at runtime.
@@
Max-Info Path Planning: Given a partially-known environment and metrics for picture evaluation, generate a path that minimizes the time to the inspecting task while ensuring the robot’s safety. The path planner should also accommodate changes in the environment, as new obstacles may appear as the vehicle approaches the best inspection point.
@@ IV Methodology
The proposed framework aims to provide a comprehensive strategy to inspect a target with the fewest number of viewpoints. In defining such an inspection viewpoint, the target object should occupy a certain proportion of the frame from the viewpoint’s perspective, minimizing distortion and reducing any obstructions. If the target is not fully visible in one frame, the robot must decide autonomously to focus on one part of the target and subsequently move to capture the remainder of the target. In this work, we propose an objective function that mathematically defines what a quality inspection photo is and a method to capture quality images of the entire target with as few pictures as possible. Referring again to Fig. 1, we point out that the house is best viewed from viewpoint D since it is minimally obscured and distorted, while other views show little of the house or have a distorted view of the target.
Fig. 2 shows an overview of our approach to guide the autonomous mobile robot during the mission. As shown, the vehicle is initialized at a starting point in a partially-known environment with the target information known. The vehicle updates its map using its equipped depth sensor. After the update, candidate views are generated and evaluated via our quality scoring metrics. We use particle swarm optimization (PSO) to iteratively search for the best viewpoint in the environment but any DFO method can be utilized. The best viewpoint is then passed to the vehicle’s planner and controller to drive it toward the location. As the vehicle moves, the best viewpoint is reassessed with any new information, given repeated map updates. Once the vehicle reaches the best viewpoint, a photo is captured and the target interest is updated based on the estimated target coverage from the captured photo using Gaussian process interpolation. Our approach is discussed in detail in the following sections, starting with the photo quality metrics followed by candidate view generation and evaluation.
@@ IV-A Photo Evaluation Metric
First, we introduce the metric to determine the ideal position to capture an image of a desired target. To aid the reader’s comprehension, we begin with a visual illustration of the metric using a two-dimensional example in Fig. 3. Furthermore, we offer a similar three-dimensional example in Fig. 4 that showcases the expansion of the two-dimensional metrics to three dimensions. We consider a discretized target $\mathbf{T}\in\mathbb{R}^{m\times n_{d}}$ where $n_{d}$ is the dimension of the target and defined by $m$ coordinates. In the 2D example, we show a UAV positioned slightly behind an obstacle, obscuring its view of the full target. For the visible portion of the target $\mathbf{T}_{v}\subseteq\mathbf{T}$ and for $\tau\in\mathbf{T}_{v}$, $g(\tau,\bm{p})=1$ indicates that the visible portion has not been inspected previously and is visible from viewpoint $\bm{p}$, while $g(\tau,\bm{p})=0$ indicates that the region is obscured or has already been observed in the past. In our approach, we seek to autonomously navigate and find the optimal location or a few locations to fully inspect the desired object. We define the optimal viewpoint $\bm{p}^{*}$ as follows:
@@
$\textrm{where}\;\;G(\mathbf{T},\mathcal{M},\bm{p})=\gamma_{d}\cdot\gamma_{s}\cdot\sum_{\tau\in\mathbf{T}}g(\tau,\bm{p}).$ (2)
The function $g(\tau,\bm{p})$ signifies the level of interest that remains for any portion of the target after viewing from viewpoint $\bm{p}$. As shown in Fig. 3, the visible portion of the target is a function of the position $\bm{p}$ and the occupancy map $\mathcal{M}$. The variables $\gamma_{d}$ and $\gamma_{s}$ are discount factors associated with utility functions $U_{d}$ and $U_{s}$ to assess the quality of a viewpoint given its respective camera parameters, position, and environmental obstructions. These factors are essential in determining the robot’s choice of optimal viewpoints, as they influence the amount of distortion that can be accepted in the inspection pictures and the amount of the target that should occupy the image. The evaluation of these factors will be discussed in more detail in the following sections.
@@ IV-A1 Perspective Distortion Discount Factor
Directly facing an inspected surface is often the ideal view for inspection tasks to maximize the amount of information in one photograph and create a consistent view of the target [15]. In fields such as product photography, architecture, or real estate where precise measurements and representations are required, if the view is at an angle, structures can appear distorted, which may compromise the resulting image capture. We introduce a perspective distortion discount factor that penalizes views that cause a distortion of the target. The target distortion in our approach is formally defined as follows:
@@
where $\ell_{i1}$ and $\ell_{i2}$ represent the maximum distances from the optical center associated with the intersection of the camera frame and the target. The variable $\ell_{i\mathbf{T}}$ is the length of the target captured along dimension $i$. $U_{d}$ represents the user-defined utility function associated with the perspective distortion, penalizing unbalanced target captures, and $\mathbf{q}_{d}$ is a vector of parameters for the user-defined function. For example, in a 2D scenario if $\ell_{11}=\ell_{12}$, the view of the vehicle’s camera is perpendicular to the target, minimizing perspective distortion and maximizing $\gamma_{d}$. If $\ell_{11}>>\ell_{12}$ or $\ell_{11}<<\ell_{12}$, the vehicle is positioned far to the left or right of the target, respectively, maximizing the perspective distortion and resulting in a small $\gamma_{d}$. This factor is designed to account for information loss that may occur in side views due to the inherent limitations of perspective projection. By incorporating this factor into the optimization process, the proposed method can effectively balance the trade-off between the ideal view and the available view, improving the overall quality of inspection images.
@@ IV-A2 Scale Discount Factor
The term scale in photography refers to the relative size of the target versus all other objects in the frame. We define a favorable picture as a picture in which the target is in the foreground and the target occupies a percentage $\beta$ of the frame. Accurate sizing of a subject within the frame is essential for an accurate representation of the target. If the scaling is incorrect, the image quality will be poor and there will be little useful information. An example of this is in real estate applications, where a properly sized image will give prospective buyers an idea of the layout of the property and the design elements of the home. We use $\gamma_{s}$ to represent the scale discount factor that encourages an inspection image with a percentage of the frame being the target of interest. The penalty factor is determined using the target’s estimated area within the frame. The variable $\beta$ is a user-defined parameter that specifies the desired optimal percentage of the target that is within the frame of the picture. By applying a discount to target coverage using the scale discount factor, we ensure that the optimal image captured by the robot contains an appropriate and desirable proportion of the target object. We use the following equation to calculate this factor:
@@
where $\phi_{i\min}$ and $\phi_{i\max}$ are the minimum and maximum angles of the intersection between the target plane and center of the frame along dimension $i$ and $\Phi_{i}$ is the field of view of the camera along dimension $i$. We denote $U_{s}$ as the user-defined utility function associated with the scale of the target in the frame and $\mathbf{q}_{s}$ as a vector of parameters for the user-defined function. In (4) , if $\phi_{i\min}=-\Phi/2$ and $\phi_{i\max}=\Phi/2$, the target would occupy $100\%$ of the frame along dimension $i$. As a result, the utility function $U_{s}(\cdot)$ would score this percentage based on the parameter $\beta$. In this case, if $\beta=1$, the value of $\gamma_{s}$ would be maximized. To illustrate this concept, Fig. 3 presents a physical representation of the $\phi$s and the corresponding scale of the target within the vehicle’s frame.

### v2

# Take Your Best Shot: Sampling-Based Planning for Autonomous Photography

@@ Abstract
Autonomous mobile robots (AMRs) equipped with high-quality cameras are revolutionizing the field of autonomous photography by delivering efficient and cost-effective methods for capturing dynamic visual content. As AMRs are deployed in increasingly diverse environments, the challenge of consistently producing high-quality photographic content remains. Traditional approaches often involve AMRs following a predetermined path while capturing data-intensive imagery, which can be suboptimal, especially in environments with limited connectivity or physical obstructions. These drawbacks necessitate intelligent decision-making to pinpoint optimal vantage points for image capture. Inspired by Next Best View studies, we propose a novel autonomous photography framework that enhances image quality and minimizes the number of photos needed. This framework incorporates a proposed evaluation metric that leverages ray-tracing and Gaussian process interpolation, enabling the assessment of potential visual information from the target in partially known environments. A derivative-free optimization (DFO) method is then proposed to sample candidate views and identify the optimal viewpoint. The effectiveness of our approach is demonstrated by comparing it with existing methods and further validated through simulations and experiments with various vehicles.
Note—Code and videos of the simulations and experiments are provided in the supplementary material and can be accessed at https://www.bezzorobotics.com/sg-lb-icra25.
@@ I Introduction
Autonomous mobile robots (AMRs) are becoming increasingly prevalent in industries like manufacturing, healthcare, logistics, real-estate, and entertainment. One of the most notable applications in which AMRs play a key role is robotic photography [1, 2] due to their ability to reach view points that would otherwise be inaccessible or even hazardous for humans, e.g., disaster sites, high-rise buildings, or underwater structures [3] or simply to perform tasks more quickly or strategically [4]. For example, Fig. 1 depicts a UAV tasked to take the most comprehensive picture of a house. As can be noted by the different viewpoint pictures, some angles are better than others due to the presence of obstacles that occlude the target object of interest. This ability to assess and select the best viewpoint (i.e., D in the figure) for capturing critical information is essential to the success of robotic photography tasks, especially in dynamic and cluttered environments.
Despite these advancements, one of the key challenges in automating robotic photography is determining what constitutes a “good” picture. The evaluation of photo quality is inherently subjective, and even human observers often disagree on what makes an image satisfactory [5, 6]. This subjectivity complicates the creation of a formal metric for photo quality assessment. Additionally, current robotic systems often collect large amounts of redundant or irrelevant data, requiring extensive human post-processing to sort and evaluate the captured images.
In this work, we identify two significant gaps in the existing literature: i) the absence of a clear, formal definition of a high-quality picture in the context of robotic photography, and ii) the lack of efficient data collection frameworks that reduce unnecessary images and minimize human involvement in post-processing. Addressing these gaps is crucial to making autonomous photography more practical and scalable in real-world operations. By developing better tools for assessing image quality and optimizing data collection, robots will be able to complete tasks more autonomously, reducing the need for human intervention and increasing overall efficiency.
@@
In our autonomous photography framework, we propose an evaluation metric based on perspective distortion, the scale of a target within the viewing frame, and the estimated target coverage which will allow us to find the best viewpoint. The metric is used in conjunction with a derivative-free optimization (DFO) method which samples the environment to find the best viewpoint. The main contribution of this work is the development of an autonomous photography framework that utilizes computationally efficient Gaussian process interpolation and derivative-free optimization (DFO) to optimize our proposed evaluation metric over uniformly sampled candidate views. This framework enables the runtime capture of high-quality and aesthetically pleasing images of the target in a partially known environment.
Figure 1: A UAV captures a photograph of a house. Angles A, B, and C represent obstructed or limited views, while a bird’s-eye view (D) offers the most comprehensive information with minimal obstruction and distortion.
@@ II Related Work
This paper addresses a unique challenge in robotic photography, distinct from autonomous inspection. Although both fields aim to capture insightful images, autonomous inspection focuses on detailed visual documentation for precise target reconstruction or digital replication. In contrast, the photographic approach prioritizes obtaining an overall view of the subject. Our methodology considers that some parts of the target may be missed due to obstructions or occlusions, situations that are less acceptable in inspection contexts where every detail is crucial. Instead, the approach prioritizes capturing the overall essence of the target, similar to human photographers, aiming to gather as much information as possible with few observations.
Next-Best-View (NBV) algorithms, first introduced by Connolly in [7], focus on the exploration of objects or environments by selecting the most informative next viewpoint for a robot, optimizing this choice based on the robot’s goals and constraints. NBV is especially important in 3D object reconstruction, where determining the optimal next viewpoint is crucial. In more recent studies, Naazare et al. [8] proposed a weighted multi-objective optimization approach to select NBVs for a mobile robotic arm, while Han et al. [9] used a double-branch network architecture to rank NBVs. Dhami et al. [10] extended these ideas by employing two robotic arms for a more efficient reconstruction. However, despite the success of these methods in acquiring dense data, they tend to prioritize collecting information near the target without fully addressing data redundancy or the appearance of the target in the captured photograph.
Another challenge of directly applying NBV-based solutions to autonomous photography is computing the optimal viewpoint in partially known or unknown environments, where unexpected objects can not only obscure the target but also pose safety threats to vehicles. Authors in [11] navigate robots to optimal viewpoints in known environments, while the authors in [12] investigate a receding horizon NBV, utilizing a random tree method to guide the robot along a path in an unknown environment. [13] proposes a guided NBV approach for large-scale 3D reconstruction, which requires a rough global scan prior to a detailed NBV inspection. [14] computes NBV based on the map built during frontier exploration. In contrast, our work focuses on applications where the target location is known but the surrounding environment is unknown such as real-estate or surveillance. Given the target location, the robot dynamically updates the optimal viewpoint to efficiently adapt to changes in the environment.
@@ III Problem Formulation
For typical photography missions, target information is known, but the surroundings can be uncertain. We assume the robot knows the target’s location and dimensions $\mathbf{T}$. Obstacles, represented as an occupancy map $\mathcal{M}$, can be known or unknown $\mathcal{O}$. The vehicle should update its viewpoint when detecting obstacles that block the target. Equipped with range sensors (LiDAR, sonar, cameras), the robot can recognize obstacles and measure its distance to the target. The goal is to maximize the information captured in the photo while reducing the number of pictures needed during operations. To achieve this, we formally define the problems as follows:
@@
Metrics of the Best View for Capturing Target: Consider the location and dimensions of the target represented by the coordinate set $\mathbf{T}$ and a set of initially unknown obstacles that are discovered at runtime and represented by an occupancy map $\mathcal{M}$. The goal is to find an evaluation metric $G(\mathbf{T},\mathcal{M},\mathbf{P})$ that scores the sampled viewpoints $\mathbf{P}$ for the vehicle to visit at runtime.
@@
Max-Info Path Planning: Given a partially-knownn environment and metrics for picture evaluation, generate a path that minimizes the time to the viewpoint while ensuring the robot’s safety. The path planner should also accommodate changes in the environment, as new obstacles may appear as the vehicle approaches the best viewpoint.
@@ IV Methodology
The framework aims to autonomously capture the entire target from as few viewpoints as possible. The camera position ensures the target occupies a specific proportion of the frame, minimizing distortion and obstructions. If the target is not fully visible in one frame, the robot must decide autonomously to focus on one part of the target and subsequently move to capture the remainder of the target. In this work, we propose an objective function that formally defines the quality of a photograph and a method to capture quality photographs of the entire target with as few pictures as possible.
Fig. 2 describes a method for guiding an autonomous mobile robot during a mission in a partially-known environment. Starting from a known point, the robot utilizes a depth sensor to update its map and interest in unseen portions of the target while generating candidate views. These views are scored for quality, and particle swarm optimization (PSO) is employed to determine the optimal camera position, though any DFO method could be applied. The robot continuously re-evaluates its viewpoint as it moves, updating its target coverage using Gaussian process interpolation once the optimal point is reached and a photo is taken. Further details on metrics and view evaluation follow.
@@ IV-A Photo Evaluation Metric
First, we introduce the metric to determine the ideal position to capture an image of a desired target which were chosen based on careful discussion with our research sponsors in the real-estate industry. To aid the reader’s comprehension, we begin with a visual illustration of the metric using a two-dimensional example in Fig. 3 followed by an example in Fig. 4 that showcases the expansion of the 2D metrics to 3D. We consider a discretized target $\mathbf{T}\in\mathbb{R}^{m\times n_{d}}$ where $n_{d}$ is the dimension of the target and defined by $m$ coordinates. In Fig. 3, a UAV is positioned slightly behind an obstacle, obscuring its view of the full target. For the visible portion of the target $\mathbf{T}_{v}\subseteq\mathbf{T}$ and for $\tau\in\mathbf{T}_{v}$, $g(\tau,\bm{p})=1$ indicates that the visible portion has not been captured previously and is visible from viewpoint $\bm{p}$, while $g(\tau,\bm{p})=0$ indicates that the region is obscured or has already been observed in the past, making the interest in that section low or unimportant to recapture. We define the optimal viewpoint $\bm{p}^{*}$ as follows:
@@
$\textrm{where}\;\;G(\mathbf{T},\mathcal{M},\bm{p})=\gamma_{d}\cdot\gamma_{s}\cdot\sum_{\tau\in\mathbf{T}}g(\tau,\bm{p}).$
The function $g(\tau,\bm{p})$ signifies the level of interest that remains for any portion of the target after viewing from viewpoint $\bm{p}$. As shown in Fig. 3, the visible portion of the target is a function of the position $\bm{p}$ and the occupancy map $\mathcal{M}$. The variables $\gamma_{d}$ and $\gamma_{s}$ are discount factors associated with utility functions $U_{d}$ and $U_{s}$ to assess the quality of a viewpoint given its respective camera parameters, position, and environmental obstructions. These factors are essential in determining the robot’s choice of optimal viewpoints, as they influence the amount of distortion that can be accepted in the pictures and the amount of the target that should occupy the image. The evaluation of these factors will be discussed in more detail in the following sections.
@@ IV-A1 Perspective Distortion Discount Factor
Directly facing the target surface is often the ideal view to maximize the amount of information in one photograph and create a consistent view of the target [15]. In fields such as product photography, architecture, or real estate where precise measurements and representations are required, distorted structure is less preferred. We introduce a perspective distortion discount factor that penalizes views that cause a distortion of the target. The target distortion in our approach is formally defined as follows:
@@
where $\ell_{i1}$ and $\ell_{i2}$ represent the maximum distances from the optical center associated with the intersection of the camera frame and the target. The variable $\ell_{i\mathbf{T}}$ is the length of the target captured along dimension $i$. $U_{d}$ represents the user-defined utility function associated with the perspective distortion, penalizing unbalanced target captures, and $\mathbf{q}_{d}$ is a vector of parameters for the user-defined function. For example, in a 2D scenario if $\ell_{11}=\ell_{12}$, the view of the vehicle’s camera is perpendicular to the target, minimizing perspective distortion and maximizing $\gamma_{d}$. If $\ell_{11}>>\ell_{12}$ or $\ell_{11}<<\ell_{12}$, the vehicle is positioned far to the left or right of the target, respectively, maximizing the perspective distortion and resulting in a small $\gamma_{d}$. By incorporating this factor into the optimization process, the proposed method can effectively balance the trade-off between the ideal view and the available view, improving the photo quality.
@@ IV-A2 Scale Discount Factor
In photography, the term scale refers to the relative size of the target compared to other objects in the frame. A favorable picture places the target in the foreground, occupying a percentage $\beta$ of the frame. Proper scaling is crucial for accurately representing the target. The scale discount factor $\gamma_{s}$ is used to encourage capturing images where the target occupies the desired percentage of the frame, determined by $\beta$. We use the following equation to calculate this factor:
@@
where $\phi_{i\min}$ and $\phi_{i\max}$ are the minimum and maximum angles of the intersection between the target plane and center of the frame along dimension $i$ and $\Phi_{i}$ is the field of view of the camera along dimension $i$. We denote $U_{s}$ as the user-defined utility function associated with the scale of the target in the frame and $\mathbf{q}_{s}$ as a vector of parameters for the user-defined function. In (3) , if $\phi_{i\min}=-\Phi_{i}/2$ and $\phi_{i\max}=\Phi_{i}/2$, the target would occupy $100\%$ of the frame along dimension $i$. The utility function $U_{s}(\cdot)$ would score this percentage based on how close it is to the desired percentage $\beta$. Fig. 3 and Fig. 4 present a physical representation to illustrate this concept.